tan15.tan25.tan35=tanx
find x in degree where x belongs to(0,45)

Grade:12

1 Answers

Aman Bansal
11 years ago

Dear Anil

1. Using the identity: tan x * tan(60° - x) * tan(60° + x) = tan(3x) well have for x = 5°:
tan 5° tan 55° tan 65° = tan 15° = cot 75° = 1/tan 75°, hence
tan 5° tan 55° tan 65° tan 75° = 1 or
tan 5° = 1/(tan 55° tan 65° tan 75°) = cot 55° cot 65° cot 75° =
= tan 35° tan 25° tan 15°, so the answer is t = 5°
Proof of the identity: tan x * tan(60° - x) * tan(60° + x) =
(sin x * sin(60° - x) * sin(60° + x)) / (cos x * cos(60° - x) * cos(60° + x)) =
= (sin x * (cos 2x + 1/2)) / (cos x * (cos 2x - 1/2)) =
= ((1/2)(sin 3x - sin x) + (1/2) sin x) / ((1/2)(cos 3x + cos x) - (1/2) cos x) = tan 3x

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert